OpenStudy (anonymous):

what is the simplest form of 16 divided by the square root of 6

OpenStudy (ipwnbunnies):

\[\frac{ 16 }{ \sqrt{6}}\]

OpenStudy (ipwnbunnies):

That? Unless evaluated, that's the simplest form I can see.

OpenStudy (ipwnbunnies):

@magicstudy1903

Parth (parthkohli):

Of course the word "simple" is subjective, but people hate irrational numbers in denominators. So maybe that's what we have to simplify.

OpenStudy (ipwnbunnies):

Ahh, you're right.

Parth (parthkohli):

To rationalize the denominator, multiply both numerator and denominator by \(\sqrt 6\)